How do i find the exact values of x for which 3(2x-1)^2+2(2x-1)-8=0?

Answer by lwsshak3(11628)

How do i find the exact values of x for which
3(2x-1)^2+2(2x-1)-8=0?
..
let u=(2x-1)
u^2=(2x-1)^2
..
3u^2+2u-8=0
(3u-4)(u+2)=0
u=4/3=2x-1
6x-3=4
6x=7
x=7/6
or
u=-2=2x-1
2x=-1
x=-1/2